OpenStudy (biue):

Which is equal to 6 - 2?

OpenStudy (anonymous):

2+2

OpenStudy (biue):

The choices are 7 - 1 , 2 + 4 and 3 + 1

OpenStudy (lynfran):

6-2=4 now which option adds up to that?

OpenStudy (biue):

2 + 4?

OpenStudy (lynfran):

2+4=6 nope

OpenStudy (biue):

Crap hmm

OpenStudy (biue):

7 - 1?

OpenStudy (lynfran):

7-1=6 nope

OpenStudy (biue):

so 3 + 1?

OpenStudy (biue):

3+1=4

OpenStudy (lynfran):

correct

OpenStudy (biue):

Thank you for helping me

OpenStudy (lynfran):

welcome